OpenStudy (anonymous):

\[\sqrt{b^2-4}=b-2\] symbolically, im not liking these lol

OpenStudy (anonymous):

square both isdes and get \[b^2-4=(b-2)^2\] \[b^2-4=b^2-4b+4\] \[-4=-4b+4\] \[-8=-4b\] \[b=2\] and we check

OpenStudy (anonymous):

have to check because we squared. replace x by 2, get zero on both sides so it is right

OpenStudy (anonymous):

how do you do these so fast? it is like pulling teeth with me

OpenStudy (anonymous):

you have a square root, so what choice do you have but to square? the rest is algebra practice

OpenStudy (anonymous):

which is my weak spot